OpenStudy (anonymous):

Can someone help me solve this step by step 2x+12/(x-2)=x/(x-2)+5

OpenStudy (anonymous):

\[2x+\frac{ 12 }{ x-2}= \frac{ x }{ x-2 }+5\]

OpenStudy (anonymous):

1- multiply both side by (x-2) 2- move x in one side and number to the other side

OpenStudy (anonymous):

2x(x-2)+12=x+5(x-2) 2x^2-4x+12=5x-10 2x^2-4x-5x=-10-12 2x^2-9x+22=0